OpenStudy (anonymous):

lim as x approaches 0 of (-2)/(cos(x))

OpenStudy (anonymous):

what do you get when you replace \(x\) by \(0\) ?

OpenStudy (anonymous):

it would be -2 but sorry i meant to say as x approaches pi/2

OpenStudy (anonymous):

what do you get when \(x\) approaches \(\frac{\pi}{2}\) ?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

let me rephrase that what you you get literally when \(x=\frac{\pi}{2}\) ? hint, it is certainly not zero

OpenStudy (anonymous):

so -2/cos(pi/2) would be DNE?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you get \[-\frac{2}{0}\] so there is no limit

OpenStudy (anonymous):

if you got \[\frac{0}{0}\] you would have more work to do

OpenStudy (anonymous):

oh ok thank you and one more, how do i do limit as x approaches 9pi/4 of( cos(x)-1)/(6x)? i got -sqrt(2)-1/(27pi) but i dont think thats correct